Find the angle between the curves y^2=ax and ay^2=x^3 (a>0), at the points of intersection other than origin.


1
Expert's answer
2021-05-25T15:37:54-0400

"\text{Given, the two curves} 
y ^2 = ax \space and \space
ay^2 = x ^3.\newline
\text{Equating the both equations}
y ^2 = ax \space
and \space
ay^2 = x ^3 \text{to get the integrating point,}\newline
ax = \frac{x^3}{a}\newline
x ^2 = a ^2\newline
x = a\newline
 \text{Then,}\newline
y = (a^2) ^{\frac{1}{2 }}= a.\newline 
\text{The integrating point is }
 \space
(a,a).\newline 
\text{Differentiate both curves w.r.t} 
 \space
x.\newline 
\text{First curve}\newline 
y ^2 = ax\newline
2y \frac{dy}{dx}= a\newline
 \frac{dy}{dx}= \frac{a}{2y}\newline \text{Then, slope of the first curve at }
 \space
(a,a) \space
is \space
 \frac{1}{2}i.e.,m1 =\frac{1}{2}.\newline
 \text{Now, second curve,}\newline
ay^2 = x ^3\newline
2ay\frac{dy}{dx}=3x^2\newline
 \frac{dy}{dx}= \frac{3}{2}\newline \text{Therefore, } \space
m2=\frac{3}{2}.\newline
 \text{Then, the angle between the curves given by,}\newline
tan\theta= | \frac{m1 – m2}{1 + m1*m2} |\newline
 = | \frac{\frac{1}{2} – \frac{3}{2}
}{1 + \frac{1}{2}×\frac{3}{2}}|\newline
 = |\frac{-4}{7}| \newline
=\frac{4}{7}\newline
\theta=tan^{-1}(\frac{4}{7})\newline
=29.7°\newline
 \text{Thus, angle is} \space
 29.7°."